# Oryx MLLM: On-Demand Spatial-Temporal Understanding at Arbitrary Resolution

### v2

@@ 1 Introduction
Multi-Modal Large Language Models (MLLMs) have made significant strides in processing and integrating visual and linguistic inputs to generate coherent and contextually relevant responses. Proprietary models such as (OpenAI, 2023b; 2024; GeminiTeam, 2024) exemplify the cutting-edge capabilities of MLLMs. Concurrently, the open-source community is actively advancing MLLMs by enhancing their ability to understand diverse visual content (Tong et al., 2024; Liu et al., 2024g; Yang et al., 2023a), including images (Li et al., 2024a; Chen et al., 2024b), videos (Lin et al., 2023a; Cheng et al., 2024; Qian et al., 2024), and 3D data (Hong et al., 2023), etc. As MLLMs become stronger, there is a growing need for more general and unified MLLMs that are capable of processing visual content in more diverse forms and accomplishing more challenging multimodal problems.
@@ 3 Methods
In this section, we provide a detailed explanation of Oryx’s contribution. Our design is segmented into two primary components: the architecture and the training pipeline, which are elaborated upon in Section 3.1 and 3.2, respectively. We describe our innovative architecture to process native and on-demand visual inputs within MLLMs, as illustrated in Figure 2, enabling the development of a model capable of generalizing across image, video, and 3D data. Furthermore, we outline the simple yet effective training pipeline of the Oryx model.
@@ 3.1.1 Visual Representations with Native Resolution
To address the issue of native resolution processing, we introduce a visual encoder named OryxViT, which builds upon the advanced SigLIP (Zhai et al., 2023) models and is based on the Vision Transformer (Dosovitskiy, 2020) architecture. We modify the vision encoder by incorporating a sufficiently large position embedding matrix $P$ that accommodates the maximum target input sizes (2048$\times$2028 in our models, which can also be further interpolated for even larger inputs). For each visual input, we rescale the original position embeddings into $P^{\in\lfloor H/p\rfloor\times\lfloor W/p\rfloor}$ using bilinear interpolation and apply the transformation $x=x+P$. The adaptation strategy for the newly defined $P$ under native input resolution follows the training format of common MLLMs. We employ a relatively lightweight LLM as the language interface, keeping the vision encoder’s parameters unfrozen while freezing most of the other parameters. We collect training data pairs from multiple vision-language tasks including captioning, OCR, visual question answering, etc.
A significant challenge is managing the dynamic sequence length $N=\lfloor H/p\rfloor\times\lfloor W/p\rfloor$ for the Vision Transformer during batch processing, where we propose the Variable-Length Self-Attention strategy to address this issue. For visual patches with lengths $N_{1},N_{2},\ldots,N_{b}$ in a batch of size $b$, we concatenate the patches across the sequence dimensions into a shape of $[1,\sum_{i=1}^{b}N_{i},C]$ before feeding them into the transformer blocks. We utilize the variable-length attention operator provided in flash attention (Dao et al., 2022) to compute the attention for each visual input within the batch independently. With these designs, our OryxViT can efficiently process visual signals of varying aspect ratios in batch mode, maintaining a forward speed comparable to that of conventional fixed-resolution visual encoders.
@@ 3.1.2 On-Demand Dynamic Compression Supporting Long Visual Context
Using the visual representation feature map $f$, the compression serves as the bridge between vision and language modalities. We implement downsample layers with varying ratios to accommodate different input lengths. Specifically, we categorize the visual context into pure images, short videos, and long videos, applying downsample layers $d_{1},d_{2},d_{3}$ respectively. In our implementation, we set $d_{3}=4d_{2}=16d_{1}$, therefore the token length of frames in the long videos is reduced to $\frac{1}{16}$ of that of images with the same resolution.
@@ 3.1.3 One Model for All: Image, Video, and 3D Understanding
Long-Form Temporal Training with Needle-In-A-Haystack. The key ability for processing long-form video inputs is the identification of specific information within an extensive context, akin to the ”needle-in-a-haystack” task in the NLP field. To enhance the Oryx model’s capability to pinpoint details, we prepare long-form temporal needle-in-a-haystack training data. Specifically, we source video samples from the MovieNet (Huang et al., 2020) dataset, which comprises an average of 1000 frames per movie and an average duration of 45 minutes, thereby providing a natural setting for retrieving designated targets. We devise two tasks to train the model: captioning and differing. The captioning task requires the model to generate captions for frames at specific indices, while the differing task involves identifying differences between two frames given their indices. The training corpus is generated using GPT-4o, which produces captions for single frames or frame pairs. These captioned frames are then reinserted into the overall movie sequences, ensuring the training data maintains contextual integrity.
@@
3.2 Training Pipeline & Data Mixture
@@
Stage 1: Text-Image Pre-training and Supervised Fine-tuning. In the first stage of our training process, we focus on developing the foundational vision-language capabilities of the Oryx model using image data. This stage begins with a pre-training phase to train the dynamic compressor component with a dataset of 558k images in LLaVA-1.5 (Liu et al., 2024b). Following this, we gather a collection of 4 million supervised fine-tuning image-text pairs that focus on high-quality knowledge learning. This data is sourced from various open-source academic datasets, including LLaVA-NeXt (Liu et al., 2024c), Cauldron (Laurençon et al., 2024), and Cambrian-1 (Tong et al., 2024). It is important to note that we do not incorporate large-scale pre-training stages as described in (Li et al., 2024a) or employ exclusive supervised fine-tuning data such as those in (Lin et al., 2023b; Bai et al., 2023), as our primary objective is to validate the effectiveness of our unified Oryx architecture.
Stage 2: Joint Supervised Fine-tuning. In Stage 2, we further conduct a supervised fine-tuning procedure following the initial stage, aiming to jointly train the Oryx model with image, video, and 3D-aware visual inputs. The image training data is sampled from the dataset collected during the supervised fine-tuning phase of Stage 1, ensuring a balanced ratio of image and video data by utilizing around 600k image-text pairs. For video data, we source both comprehensive and multiple-choice datasets from open-source video repositories. Comprehensive datasets, which include question-answering and captioning tasks, are integrated using VideoChatGPT-Plus (Maaz et al., 2024), ShareGPT4Video (Chen et al., 2024a) and LLaVA-Hound (Zhang et al., 2024b). To enhance performance on multiple-choice benchmarks, we further incorporated Cinepile (Rawal et al., 2024), NextQA (Xiao et al., 2021) and PerceptionTest (Patraucean et al., 2024) into our training dataset. Additionally, we include video samples of needle-in-a-haystack data generated by GPT-4o (OpenAI, 2024) for long-form video learning and spatial-aware 3D multi-frame samples from the ScanQA (Azuma et al., 2022) training dataset, culminating in a total of around 650k video samples. The supervised fine-tuning strategy in this stage mirrored that of Stage 1, ensuring consistency in the training approach.
@@ Appendix A Generation Results
Video Summarization and Detailed Description. As shown in Fig. 5, the Oryx model effectively generates a comprehensive and detailed caption that accurately summarizes the input video. It captures the main event while preserving essential information.
@@
Figure 6: Oryx learns to reason through the input video.
Skill Learning From Videos. Oryx can acquire useful skills from the input video. As demonstrated in Fig. 7, Oryx learns to use Google Scholar to cite a paper by following the steps shown in the video. It illustrates all the necessary steps to complete the citation, highlighting its strong skill-learning capability and potential for agent-based tasks and task execution.
@@
Understanding 3D with Coarse Correspondences. Oryx enhances its 3D spatial understanding using coarse correspondences. Fig. 8 illustrates Oryx’s reasoning process, demonstrating its ability to improve 3D comprehension through these correspondences and generate accurate reasoning outcomes.
@@
Appendix B Training Details

### v3

@@ 1 Introduction
Multi-Modal Large Language Models (MLLMs) have made significant strides in processing and integrating visual and linguistic inputs to generate coherent and contextually relevant responses. Proprietary models such as (OpenAI, 2023b; 2024; GeminiTeam, 2024) exemplify the cutting-edge capabilities of MLLMs. Concurrently, the open-source community is actively advancing MLLMs by enhancing their ability to understand diverse visual content (Tong et al., 2024; Liu et al., 2024g; Yang et al., 2023a; Dong et al., 2024), including images (Li et al., 2024a; Chen et al., 2024b), videos (Lin et al., 2023a; Cheng et al., 2024; Qian et al., 2024), and 3D data (Hong et al., 2023), etc. As MLLMs become stronger, there is a growing need for more general and unified MLLMs that are capable of processing visual content in more diverse forms and accomplishing more challenging multimodal problems.
@@ 3 Methods
In this section, we provide a detailed explanation of Oryx’s contribution. Our design is segmented into two primary components: the architecture and the data curation & training pipeline, which are elaborated upon in Section 3.1 and 3.2, respectively. We describe our innovative architecture to process native and on-demand visual inputs within MLLMs, as illustrated in Figure 2, enabling the development of a model capable of generalizing across image, video, and 3D data. Furthermore, we outline the simple yet effective training pipeline of the Oryx model.
@@ 3.1.1 Visual Representations with Native Resolution
To address the issue of native resolution processing, we introduce a visual encoder named OryxViT, which builds upon the advanced SigLIP (Zhai et al., 2023) models and is based on the Vision Transformer (Dosovitskiy, 2020) architecture. We modify the vision encoder by incorporating a sufficiently large position embedding matrix $P$ that accommodates the maximum target input sizes. For each visual input, we rescale the original position embeddings into $P^{\in\lfloor H/p\rfloor\times\lfloor W/p\rfloor}$ using bilinear interpolation and apply the transformation $x=x+P$. The pre-training strategy for the proposed visual encoder OryxViT under native input resolution follows the training format of common MLLMs. We employ a relatively lightweight LLM as the language interface, keeping the vision encoder’s parameters unfrozen while freezing most of the other parameters. Details for the training settings and datasets can be referred to in the Appendix.
A significant challenge is managing the dynamic sequence length $N=\lfloor H/p\rfloor\times\lfloor W/p\rfloor$ for the Vision Transformer during batch processing, where we propose the Variable-Length Self-Attention strategy to address this issue. For visual patches with lengths $N_{1},N_{2},\ldots,N_{b}$ in a batch of size $b$, we concatenate the patches across the sequence dimensions into a shape of $[1,\sum_{i=1}^{b}N_{i},C]$ before feeding them into the transformer blocks. We utilize the variable-length attention operator provided in flash attention (Dao et al., 2022) to compute the attention for each visual input within the batch independently. With these designs, our OryxViT can efficiently process visual signals of varying aspect ratios in batch mode, maintaining a forward speed comparable to that of conventional fixed-resolution visual encoders. We also provide the detailed speed test for OryxViT in the Appendix to demonstrate the efficiency.
@@ 3.1.2 On-Demand Dynamic Compression Supporting Long Visual Context
Using the visual representation feature map $f$, the compression serves as the bridge between vision and language modalities. We implement downsample layers with varying ratios to accommodate different input lengths. Specifically, we categorize the visual context into pure images, short videos, and long videos, applying downsample layers $d_{1},d_{2},d_{3}$ respectively. We satisfy the downsampling ratio $r_{1}<r_{2}<r_{3}$ for layer $d_{1},d_{2},d_{3}$ to reduce the token length for frames in videos.
@@ 3.2.1 One Model for All: Image, Video, and 3D Understanding
Long-Form Temporal Training with Needle-In-A-Haystack. The key ability for processing long-form video inputs is the identification of specific information within an extensive context, akin to the ”needle-in-a-haystack” task in the NLP field. To enhance the Oryx model’s capability to pinpoint details, we prepare long-form temporal needle-in-a-haystack training data. Specifically, we source video samples from the MovieNet (Huang et al., 2020) dataset, which comprises an average of 1000 frames per movie and an average duration of 45 minutes, thereby providing a natural setting for retrieving designated targets. We devise two tasks to train the model: captioning and differing. The captioning task requires the model to generate captions for frames at specific indices, while the differing task involves identifying differences between two frames given their indices. The training corpus is generated using SOTA LLMs, which produces captions for single frames or frame pairs. These captioned frames are then reinserted into the overall movie sequences, ensuring the training data maintains contextual integrity.
@@
3.2.2 Training Pipeline & Data Mixture
@@
Stage 1: Text-Image Pre-training and Supervised Fine-tuning. In the first stage of our training process, we focus on developing the foundational vision-language capabilities of the Oryx model using image data. This stage begins with a pre-training phase to train the dynamic compressor component with the basic image captioning data. Following this, we gather a collection of 4 million supervised fine-tuning image-text pairs that focus on high-quality knowledge learning. This data is sourced from various open-source academic datasets, including LLaVA-NeXt (Liu et al., 2024c), Cauldron (Laurençon et al., 2024), and Cambrian-1 (Tong et al., 2024). It is important to note that we do not incorporate large-scale pre-training stages as described in (Li et al., 2024a) or employ exclusive supervised fine-tuning data such as those in (Lin et al., 2023b; Bai et al., 2023), as our primary objective is to validate the effectiveness of our unified Oryx architecture.
Stage 2: Joint Supervised Fine-tuning. In Stage 2, we further conduct a supervised fine-tuning procedure following the initial stage, aiming to jointly train the Oryx model with image, video, and 3D-aware visual inputs. The image training data is sampled from the dataset collected during the supervised fine-tuning phase of Stage 1, ensuring a balanced ratio of image and video data. For video data, we source both comprehensive and multiple-choice datasets from open-source video repositories. Comprehensive datasets, which include question-answering and captioning tasks, are integrated using VideoChatGPT-Plus (Maaz et al., 2024), ShareGPT4Video (Chen et al., 2024a) and LLaVA-Hound (Zhang et al., 2024b). To enhance performance on multiple-choice benchmarks, we further incorporated Cinepile (Rawal et al., 2024), NextQA (Xiao et al., 2021) and PerceptionTest (Patraucean et al., 2024) into our training dataset. Additionally, we include video samples of needle-in-a-haystack data generated by GPT-4o (OpenAI, 2024) for long-form video learning and spatial-aware 3D multi-frame samples from the ScanQA (Azuma et al., 2022) training dataset, culminating in a total of around 650k video samples. The supervised fine-tuning strategy in this stage mirrored that of Stage 1, ensuring consistency in the training approach.
@@ Appendix A Generation Results
Video Summarization and Detailed Description. As shown in Fig. 4, the Oryx model effectively generates a comprehensive and detailed caption that accurately summarizes the input video. It captures the main event while preserving essential information. Oryx produces more accurate results about the match information, the name, and the status of the player. In contrast, LLaVA-OneVision (Li et al., 2024a) shows the wrong name, and LongVILA tells the wrong score on the board.
@@
Figure 5: Oryx learns to reason through the input video.
@@
Skill Learning From Videos. Oryx can acquire useful skills from the input video. As demonstrated in Fig. 6, Oryx learns to use Google Scholar to cite a paper by following the steps shown in the video. It illustrates all the necessary steps to complete the citation, highlighting its strong skill-learning capability and potential for agent-based tasks and task execution. Although the baseline model provides additional instructions, some detailed steps are not depicted in the original video. We believe this hallucination may stem from the information in the training data.
Understanding 3D with Coarse Correspondences. Oryx enhances its 3D spatial understanding using coarse correspondences. Fig. 7 illustrates Oryx’s reasoning process, demonstrating its ability to improve 3D comprehension through these correspondences and generate accurate reasoning outcomes. In the challenging task involving direction in a first-person view, LLaVA-OneVision (Li et al., 2024a) provides incorrect conclusions.
Figure 8: Failure cases about time perception.
Figure 9: Failure cases about knowledge reasoning and transfer.
Appendix B Failure Cases
In this section, we further test the Oryx model on more challenging samples. We provide some representative failure cases to show the limitations of the Oryx model and point out the future direction for VLMs. The incorrect response is highlighted in red.
Time Perception. We provide an example of time perception in Fig. 8. From the responses, we can observe that current video understanding models using uniform frame sampling tend to lose the temporal context in videos, which is critical information between frames. For instance, Oryx provides incorrect answers and cannot accurately determine the timestamps for the given video. We believe that incorporating timestamps directly into the video or processing video inputs without frame sampling may help address this issue, leading to a more comprehensive video understanding capability.
Knowledge Reasoning and Transfer. We provide the knowledge reasoning video in Fig. 9. From the responses, we can observe that the Oryx model lacks knowledge of chemical reactions, indicating that the training data of current MLLMs is not comprehensive enough. We believe that expanding the knowledge base of current MLLMs is an urgent issue. Additionally, this example requires simple reasoning to integrate chemical knowledge into the main idea of the video, which poses a challenge for Oryx. Enhancing the reasoning capability may help address this issue.
Appendix C More Analysis
C.1 Video Needle-In-A-Haystack
C.2 Inference Speed and Efficiency
Table 6: Test on Inference Speed and Memory Cost.
Backbone Processing Approach Throughput (image/s) Max Memory Cost
OryxViT Native Resolution 146.5 49.1GB
SigLIP Dynamic Partition 157.7 48.7GB
We implement variable-length self-attention using the highly optimized FlashAttention (Dao et al., 2022) library. This allows the inference throughput of our arbitrary-resolution visual encoder to remain comparable to the dynamic partition approach used in previous methods. Additionally, the memory overhead and inference throughput remain negligible, as the variable-length attention operation is fully optimized through modifications in the CUDA kernels.
Moreover, the model size of the Vision Transformer is considerably smaller than that of large language models (400M parameters compared to 7B/32B). Consequently, the main memory cost arises from the weights and features of the LLMs, as full attention is computed on visual tokens within the LLM, even when using dynamic partitioning in visual encoding. Therefore, we maintain a similar efficiency to previous solutions in terms of inference speed and memory cost.
We tested the inference speed with an input image size of $1280\times 1280$ on one NVIDIA A100 GPU. We observed that OryxViT is only 7% slower than SigLIP with the dynamic partition approach. We believe this overhead is acceptable given the improved performance and the ability to process images at their native resolutions directly. For the max memory cost, we set the batch size to 4 and the image size to a total of $1280\times 1280$ pixels, with the aspect ratio randomly determined. The dynamic-partition baseline uses an average of 48.7G of memory, while OryxViT uses 49.1G, showing comparable results.
Figure 11: Memory-Resolution Curve for OryxViT.
Additionally, we plotted the memory-resolution curve to illustrate how the memory footprint increases with varying resolutions in Fig. 11. Our experiment was conducted on an NVIDIA A100 GPU, using square-shaped input images. The results show that the memory cost is positively correlated with image resolution and is approximately in linear relation among common image resolutions, indicating that OryxViT provides a memory-efficient solution when scaling input resolutions.
Furthermore, we can also observe that the primary memory cost arises from the model weights and the features themselves. Notably, considering that the main memory cost comes from the computation in LLMs, our method with native resolution support will not introduce much extra memory footprint compared to previous methods like LLaVA-Next (Liu et al., 2024c) and InternVL2 (Chen et al., 2024b). Oryx can largely save memory if we apply dynamic compression on visual tokens (e.g., 2x or 4x downsampling) while previous methods do not support this feature.
C.3 Ablations on Training Data
We conduct ablation experiments on the collected training data including the long-form video data and the 3D coarse corresponding data in Tab. 7. We can observe from the results that the long-form temporal data benefits the long video benchmarks. For 3D-related tasks, our proposed approach based on coarse correspondence is an effective solution for 3D benchmarks. We conducted experiments and provided results on ScanQA. ”3D Data” indicates whether 3D datasets are included in the training set, while ”C.C.” refers to using the coarse correspondence approach to annotate objects in 3D-relevant videos. We observe that both components contribute to an improved understanding results.
Table 7: Ablations on Training Data.
Long-Form Data VideoMME MLVU 55.2 58.1 55.4 59.3
3D Data C.C. METEOR ROUHE-L 11.7 28.1 12.8 32.7 14.0 35.1
C.4 Designs for MLP Adapter
Table 8: Ablations on Designs of MLP Adapter.
MLP Architecture VideoMME MLVU MMBench MMMU
Shared 55.4 59.3 81.4 43.9
Separated 54.0 54.2 81.2 43.1
We explored using specialized projection layers when integrating video modality into our model. However, we found that employing a shared MLP for all visual inputs yields better performance. To align with this design, we use the Dynamic Compressor module to closely maintain the distribution of image and video features and employ the shared MLP for visual information fusion. This ensures that input tokens for LLMs maintain a consistent distribution for both images and videos, allowing the shared MLP to be better trained with joint visual knowledge.
We conducted experiments in Tab. 8 to support our hypothesis, comparing our shared MLP strategy with separate MLPs for images and videos. Both the image and video MLPs were initialized from pre-trained image weights. The results indicate that using separate MLPs negatively impacts video benchmarks, as the dual-projector design can lead to differing distributions for similar data. Therefore, using a single MLP is a more effective solution for visual encoding.
C.5 Analysis on Downsampling
Table 9: Analysis on Downsampling.
Strategy VideoMME MLVU Average Pooling 54.6 57.5 Convolution 54.2 56.8 Q-former 42.7 35.3 Dynamic Compressor 55.4 59.3
Function VideoMME MLVU DWConv 55.0 58.9 Conv-MLP 54.7 58.5 Average Pooling 55.4 59.3
Analysis on Overall Downsampling Architecture. We integrate several mainstream approaches, including direct average pooling, $2\times 2$ spatial convolution, Q-former, and our proposed Dynamic Compressor for comparison. The results are presented in Tab. 9 (a). We reference results from VideoMME (Fu et al., 2024a) and MLVU (Zhou et al., 2024). Our observations indicate that the proposed Dynamic Compressor outperforms traditional downsampling strategies based on average pooling and spatial convolution. Additionally, we find that Q-former-based methods are not suitable for handling long visual content with fixed lengths of visual tokens. This limitation arises because the information capacity of a visual token is directly proportional to its length, making fixed lengths inadequate for more complex cases involving longer visual content.
Analysis on Downsampling Function in Dynamic Compressor We are also curious which downsampling function performs better within the dynamic compressor. We compared average pooling, DWConv, and Conv-MLP architectures. The results are presented in Tab. 9 (b). Results indicate that the parameter-free average pooling outperforms parameter-dependent methods like DWConv and Conv-MLP. This improvement is likely because average pooling better preserves the distribution of visual features, whereas more complex downsampling layers may not be effectively trained with the current pipeline.
Appendix D More Details
D.1 Implementation Details
D.2 Details of OryxViT
We pre-train OryxViT with a relatively small language model (Qwen2-0.5B (QwenTeam, 2024a) in our implementation) to enhance the language interface and improve vision-language alignment. We unfreeze OryxViT and apply LoRA fine-tuning to the language models. As a result, the total number of trainable parameters is 0.6B, making the training process significantly faster than supervised fine-tuning in the main stage (approximately 10 times faster). We collected a total of 400M pre-training data, focusing mainly on image captioning and image OCR tasks. For image captioning, we used the CapsFusion (Yu et al., 2024) datasets, and for OCR tasks, we employed synthesized OCR data pairs with OCR models. We set the batch size to 2048 and used a similar cross-entropy loss as in the main stages.
D.3 Training Details